OpenStudy (anonymous):

Write this equation in slope-intercept form y - 2.4 = 44(x-5) a. y = 2.4x + 32 b. y = 5x + 32 c. y = 2.4x - 100.6 d. y = 44x - 217.6

OpenStudy (anonymous):

which answer did u get?

OpenStudy (anonymous):

first i, pick two points on the line--for example, (2, 1) and (4, 0) . Use these points to calculate the slope: m = = = - . Next, find the y -intercept: (0, 2) . Thus, b = 2 . Therefore, the equation for this line is y = - x + 2 . Check using the point (4, 0) : 0 = - (4) + 2 yes ?

OpenStudy (anonymous):

y = mx + b y - 2.4 = 44(x - 5) y - 2.4 = 44x - 220 y -2.4 + 2.4 = 44x - 220 + 2.4 y = 44x + (- 217.6) y = 44x - 217.6 <---answer D.

OpenStudy (anonymous):

y = mx + b y - 2.4 = 44(x - 5) y - 2.4 = 44x - 220 y = 44x - 217.6 u didnt need to do the extra work :)